media%2Ff7b%2Ff7bdf477-8341-4701-831d-fePlease help
Answer #1

# CH3CH2OCH2CH3

peak at 2850 cm-1 and 2950 cm-1 for = saturated - C - H steching frequncy

peak at 1325 cm-1 for = - C - O

peak at 1100 cm-1 for = - C - C

# Ar - NH - CH3

peat at 3400 cm-1 for = N - H strching frequency

peak at 3100 cm-1 for = = C - H (aromatic)

peak at 2850 cm-1 and 2950 cm-1 for = saturated - C - H

peak at 1600 cm-1 for = C = C (aromatic)